# Positive maps and extendibility hierarchies from copositive matrices

### v1

@@ Abstract.
The characterization of positive, non-completely positive linear maps is a central problem in operator algebras and quantum information theory, where such maps serve as entanglement witnesses. This work introduces and systematically studies a new convex cone of pairwise copositive matrices, denoted $\operatorname{\mathsf{PCOP}}_{n}$. We establish that this cone is dual to the cone of pairwise completely positive matrices and, critically, provides a complete characterization for the positivity of the broad and physically relevant class of covariant maps. We provide a way to systematically lift matrices from the classical cone of copositive matrices, $\operatorname{\mathsf{COP}}_{n}$, to the new pairwise cone $\operatorname{\mathsf{PCOP}}_{n}$, thereby creating a powerful bridge between the well-studied theory of copositive forms and the structure of positive maps. We develop an analogous framework for decomposable maps, introducing the cone of pairwise decomposable matrices. As a primary application of this framework, we define a novel family of linear maps $\Phi_{t}^{G}$ parameterized by a graph $G$ and a real parameter $t$. We derive exact thresholds on $t$ that determine when these maps are positive, decomposable, or completely positive, linking these properties to fundamental graph-theoretic parameters. This construction yields vast new families of positive indecomposable maps, for which we provide explicit examples derived from infinite classes of graphs, most notably rank 3 strongly regular graphs such as Paley graphs.
On the dual side, we investigate the entanglement properties of large classes of symmetric states, such as the Dicke states. We prove that the sum-of-squares (SOS) hierarchies used in polynomial optimization to approximate the cone of copositive matrices correspond precisely to dual cones of witnesses for different levels of the PPT bosonic extendibility hierarchy. Leveraging this duality, we provide an explicit construction of bipartite (mixture of) Dicke states that are simultaneously entangled and $\mathcal{K}_{r}$-PPT bosonic extendible for any desired hierarchy level $r\geq 2$ and local dimension $n\geq 5$.
@@ 1. Introduction
The study of positive linear maps between matrix algebras is a cornerstone of modern functional analysis and operator theory [Stø63, Cho75, Wor76]. A linear map $\Phi:\mathcal{M}_{n}(\mathbb{C})\to\mathcal{M}_{n}(\mathbb{C})$ is called positive if it maps positive semidefinite matrices to positive semidefinite matrices. A stronger condition is that of complete positivity [Pau02], where the amplification $\Phi\otimes\mathrm{id}_{k}$ remains positive for all ancilla dimensions $k$. While completely positive maps admit a simple structural characterization via the Choi-Kraus representation, the set of maps that are positive but not completely positive possesses a far more intricate structure.
@@
Originating in operator algebra, the distinction between positivity and complete positivity turns out to have profound implications for quantum entanglement theory [HHHH09]. The celebrated Horodecki criterion [HHH96] established that a bipartite quantum state $\rho$ is entangled if and only if there exists a positive but not completely positive map $\Phi$ such that $\Phi\otimes\mathrm{id}](\rho)$ is not positive semidefinite. Consequently, such maps $\Phi$ are synonymous with entanglement witnesses, and the challenge of constructing and classifying them is one of the foremost problems in quantum information science [NC10, AS17, Wat18]. Within the set of positive maps lies the important convex cone of decomposable maps, that can be written as the sum of a completely positive map and a completely co-positive map (the composition of a completely positive map with the matrix transpose). Maps that are positive but not decomposable, known as indecomposable maps, are of particular interest as they can detect subtle forms of entanglement, such as that found in PPT (positive partial transpose) entangled states [Hor97]. While it was shown by Størmer and Woronowicz that every positive map $\mathcal{M}_{2}(\mathbb{C})\to\mathcal{M}_{2,3}(\mathbb{C})$ is known to be decomposable [Stø63, Wor76], Choi [Cho75] provided the first example of a positive indecomposable map $\mathcal{M}_{3}(\mathbb{C})\to\mathcal{M}_{3}(\mathbb{C})$; however, a general theory has remained largely out of reach.
[…]
$\mathsf{Sep}(\mathbb{C}^{n}\vee\mathbb{C}^{n})\subseteq\mathsf{PPTBExt}_{n}^{(\mathcal{K}_{r})}\subseteq\cdots\subseteq\mathsf{PPTBExt}_{n}^{(\mathcal{K}_{3})}\subseteq\mathsf{PPTBExt}_{n}^{(\mathcal{K}_{2})}=\mathsf{PPT}(\mathbb{C}^{n}\vee\mathbb{C}^{n}),$
and $\bigcap_{r\geq 2}\mathsf{PPTBExt}_{n}^{(\mathcal{K}_{r})}=\mathsf{Sep}(\mathbb{C}^{n}\vee\mathbb{C}^{n})$: a bosonic state is separable if and only if it is $\mathcal{K}_{r}$-PPT bosonic extendible for all $r\geq 2$. However, it was unknown for which $r$ the set $\mathsf{PPTBExt}_{n}^{(\mathcal{K}_{r})}\setminus\mathsf{Sep}(\mathbb{C}^{n}\vee\mathbb{C}^{n})$ is nonempty. Indeed, even for the minimal notion of PPT extendibility [DPS04], the explicit construction of entangled states having higher degrees of PPT extendibility has remained a challenging problem.
@@
A powerful approach to understanding cones of maps and states is through duality. The cone of positive maps is dual to the cone of separable states, while the cone of decomposable maps is dual to the cone of PPT states [EK00, SSŻ09, Kye23]. Duality also connects map properties to cones of matrices. The cone of completely positive matrices $\mathsf{CP}_{n}$ [BSM03], consisting of matrices that can be factored as $XX^{\top}$ for some entrywise non-negative (rectangular) matrix $X$, has as its dual the cone of copositive matrices $\operatorname{\mathsf{COP}}_{n}$ [SMB21]. A real symmetric matrix $A$ is in $\operatorname{\mathsf{COP}}_{n}$ if $\langle x,Ax\rangle\geq 0$ for all non-negative vectors $x\in\mathbb{R}^{n}_{+}$. Just like the set of positive maps, the cone $\operatorname{\mathsf{COP}}_{n}$ is notoriously difficult to characterize, with membership being a co-NP-complete problem [MK87]. The subset of $\operatorname{\mathsf{COP}}_{n}$ comprising matrices that are a sum of a positive semidefinite and an entrywise non-negative matrix, denoted $\mathsf{SPN}_{n}$, is tractable and corresponds to the first level of a sum-of-squares (SOS) hierarchy that approximates $\operatorname{\mathsf{COP}}_{n}$. The strict containment $\operatorname{\mathsf{COP}}_{n}\supsetneq\mathsf{SPN}_{n}$ for $n\geq 5$, exemplified by the famous Horn matrix [SMB21], mirrors the distinction between positive and decomposable maps. Let us point out here that the use of copositive matrices for entanglement detection dates back to [MATS21], see also [Mar23, RPAR+25, MMRRP+25]. On the dual side, the connection between completely positive matrices and separability dates back to [Yu16, TAQ+18], see also [JM19, SN21, Sin21, Ber23, GNS25, Gul25].
@@
A fruitful strategy for finding positive indecomposable maps is to impose symmetries on the maps under consideration. A particularly rich and physically relevant class is that of diagonal unitary covariant (DUC) maps [Liu15, SN21], which commute with conjugations by diagonal unitary matrices. Such maps, and their close relatives the conjugate diagonal unitary covariant (CDUC) maps, have recently been the subject of intense study due to their structured algebraic properties [SN22, KMP18, SDN24]. Recent work established a fundamental isomorphism between the space of DUC/CDUC maps and the vector space of pairs of matrices $(A,B)$ sharing a common diagonal [JM19, SN21]. Within this framework, a cone of pairwise completely positive matrices $\operatorname{\mathsf{PCP}}_{n}$ was introduced, which was shown to correspond precisely to the class of entanglement breaking DUC/CDUC maps. This breakthrough provided a complete description of the strongest form of positivity for these maps. However, a corresponding characterization for general positivity was left as a significant open problem. This gap motivates the central questions of this paper:
@@
    On the dual side, can these findings be applied to construct families of entangled quantum states that exhibit separability features, such as being PPT or having extensions of large order? More precisely, are there (explicit) examples of entangled bosonic states which are $\mathcal{K}_{r}$-PPT bosonic extendible?
@@
  • A bridge between the classical and pairwise cones: We establish a powerful constructive result, Theorem 4.1, which provides a canonical way to “lift” any matrix $A\in\operatorname{\mathsf{COP}}_{n}$ to a family of pairs in $\operatorname{\mathsf{PCOP}}_{n}$. This theorem enables the transfer results about copositive matrices into the domain of operator algebra and quantum information theory. These results, discussed in Section 4, and the results discussed in the next point (see Section 5) answer Q2.
  • Characterization of decomposability: We extend our analysis to decomposability by introducing the cone of pairwise decomposable matrices, $\operatorname{\mathsf{PDEC}}_{n}$. We show this cone precisely characterizes decomposable DUC/CDUC maps and prove an analogous constructive theorem (Theorem 5.6) that links $\operatorname{\mathsf{PDEC}}_{n}$ to the tractable cone $\mathsf{SPN}_{n}$. The gap between $\operatorname{\mathsf{PCOP}}_{n}$ and $\operatorname{\mathsf{PDEC}}_{n}$ thus corresponds to the domain of indecomposable positive maps, and our results bridge this gap to the one between $\operatorname{\mathsf{COP}}_{n}$ and $\mathsf{SPN}_{n}$, which has been studied extensively in the literature.
@@
  • Connections to quantum state extendibility: In the final part of the paper, we connect our results to hierarchies of quantum state extendibility. We show in Propositions 7.9 and 7.11 that the well-known sum-of-squares (SOS) hierarchies for copositive matrices ($\mathsf{K}_{n}^{(r)}$, see Section 6 and also [DKP02, Sch79]) correspond directly to witnesses for different levels of PPT-bosonic extendibility of symmetric quantum states, such as (mixture of) Dicke states. Furthermore, on the dual side, we completely characterize the PPT bosonic extendibility of Dicke states (with respect to the complete graph $\mathcal{K}_{r+2}$) in terms of the dual set $(\mathsf{K}_{n}^{(r)})^{\circ}$ (Theorem 7.15), which allows us to provide explicit examples of bosonic entangled states that are arbitrarily highly extendible whenever the local dimension $n$ is not less than $5$. This answers Q4. As a byproduct, we also prove in Corollary 7.17 the existence of $r$-partite entangled bosonic (Dicke) states of which are PPT with respect to all bipartitions.
The paper is organized as follows. Section 2 establishes our notation and reviews the necessary preliminaries on matrix cones and positive maps. Section 3 introduces the central object of study, the pairwise copositive cone $\operatorname{\mathsf{PCOP}}_{n}$, and proves its duality with $\operatorname{\mathsf{PCP}}_{n}$ as well as some other basic properties. Section 4 establishes the main constructive theorem linking the $\operatorname{\mathsf{PCOP}}_{n}$ cone to $\operatorname{\mathsf{COP}}_{n}$, hence providing a characterization of positivity for large classed of (C)DUC linear maps. Section 5 develops the analogous theory for the pairwise decomposable cone $\operatorname{\mathsf{PDEC}}_{n}$. Section 6 applies this framework to define the graph-based maps and determines the exact positivity and decomposability thresholds; we also specialize this analysis to the case of strongly regular graphs. In Section 7 we put forward the remarkable connection between hierarchies of quantum state extendibility and hierarchies of sum-of-squares programs for membership in the $\operatorname{\mathsf{COP}}_{n}$ cone. Finally, in Section 8 we discuss Markov-Choi linear maps and their positivity properties.

### v4

@@ Abstract.
The characterization of positive, non-completely positive linear maps is a central problem in operator algebras and quantum information theory, where such maps serve as entanglement witnesses. This work introduces and systematically studies a new convex cone of pairwise copositive matrices, denoted $\operatorname{\mathsf{PCOP}}_{n}$. We establish that this cone is dual to the cone of pairwise completely positive matrices and, critically, provides a complete characterization for the positivity of the broad and physically relevant class of covariant maps. We provide a way to systematically lift matrices from the classical cone of copositive matrices, $\operatorname{\mathsf{COP}}_{n}$, to the new pairwise cone $\operatorname{\mathsf{PCOP}}_{n}$, thereby creating a powerful bridge between the well-studied theory of copositive forms and the structure of positive maps. We develop an analogous framework for decomposable maps, introducing the cone $\mathsf{PDEC}_{n}$ of pairwise decomposable matrices. For several families of linear maps having diagonal unitary symmetry such as generalized Choi maps, we characterize membership in these cones using simple properties of the parameters of the maps.
As a primary application of this framework, we define a novel family of linear maps $\Phi_{t}^{G}$ parameterized by a graph $G$ and a real parameter $t$. We derive exact thresholds on $t$ that determine when these maps are positive, decomposable, or completely positive, linking these properties to fundamental graph-theoretic parameters. This construction yields vast new families of positive indecomposable maps, for which we provide explicit examples derived from infinite classes of graphs, most notably rank 3 strongly regular graphs such as Paley graphs.
On the dual side, we investigate the entanglement properties of large classes of symmetric states, such as the (mixture of) Dicke states. We prove that the sum-of-squares (SOS) hierarchies used in polynomial optimization to approximate the cone of copositive matrices correspond precisely to dual cones of witnesses for different levels of the PPT bosonic extendibility hierarchy. In the setting of the DPS hierarchy for separability, we construct a large family of boundary entanglement witnesses that are not certifiable by any level of the PPT bosonic extendibility hierarchy, answering a long standing open question from [DPS04]. Leveraging the duality, we also provide an explicit construction of bipartite (mixture of) Dicke states that are simultaneously entangled and $\mathcal{K}_{r}$-PPT bosonic extendible for any desired hierarchy level $r\geq 2$ and local dimension $n\geq 5$.
@@ 1. Introduction
The study of positive linear maps between matrix algebras is a cornerstone of modern functional analysis and operator theory [Stø63, Cho75a, Wor76]. A linear map $\Phi:\mathcal{M}_{n}(\mathbb{C})\to\mathcal{M}_{n}(\mathbb{C})$ is called positive if it maps positive semidefinite matrices to positive semidefinite matrices. A stronger condition is that of complete positivity [Pau02], where the amplification $\Phi\otimes\mathrm{id}_{k}$ remains positive for all ancilla dimensions $k$. While completely positive maps admit a simple structural characterization via the Choi-Kraus representation, the set of maps that are positive but not completely positive possesses a far more intricate structure.
@@
Originating in operator algebra, the distinction between positivity and complete positivity turns out to have profound implications for quantum entanglement theory [HHHH09]. The celebrated Horodecki criterion [HHH96] established that a bipartite quantum state $\rho$ is entangled if and only if there exists a positive but not completely positive map $\Phi$ such that $\Phi\otimes\mathrm{id}](\rho)$ is not positive semidefinite. Consequently, such maps $\Phi$ are synonymous with entanglement witnesses [Ter00], and the challenge of constructing and classifying them is one of the foremost problems in quantum information science [NC10, AS17, Wat18]. Within the set of positive maps lies the important convex cone of decomposable maps, that can be written as the sum of a completely positive map and a completely co-positive map (the composition of a completely positive map with the matrix transpose). Maps that are positive but not decomposable, known as indecomposable maps, are of particular interest as they can detect subtle forms of entanglement, such as that found in PPT (positive partial transpose) entangled states [Hor97]. While it was shown by Størmer and Woronowicz that every positive map $\mathcal{M}_{2}(\mathbb{C})\to\mathcal{M}_{2,3}(\mathbb{C})$ is known to be decomposable [Stø63, Wor76], Choi [Cho75b] provided the first example of a positive indecomposable map $\mathcal{M}_{3}(\mathbb{C})\to\mathcal{M}_{3}(\mathbb{C})$, which was later generalized in various related forms (see, e.g., [CL77, TT88, Osa91, Kye92, CKL92, Ha03, SBSC24]). However, a general theory has remained largely out of reach. Unextendible Product Bases (UPBs) underpin a range of methods for constructing positive indecomposable maps in quantum information theory. DiVincenzo et al. [DMS+03] define UPBs using orthogonality graphs, tiling patterns, and algebraic constructions; they show that the complement of a UPB yields a bound entangled state with a positive partial transpose that furnishes indecomposable maps. Terhal [Ter01] leverages these UPBs to generate positive maps that resist decomposition by standard methods. Bravyi [Bra04] classifies three‐qubit UPBs and demonstrates that uncompletable product bases yield bound entangled states, while Alon and Lovász [AL01] employ graph‐theoretic techniques to construct multipartite UPBs and derive bounds from orthogonality graphs. Several studies focus on explicit algebraic constructions of positive indecomposable maps: Chruściński and Kossakowski [CK06, CK07] introduce classes of maps based on cyclic bistochastic matrices and generalizations of the Choi map, as well as atomic maps characterized by their action on rank one projectors. Sarbicki and Chruściński [SC12] construct exposed indecomposable maps in $2n\times 2n$ matrices using reduction maps, generalizing the Robertson map. Ha and Kye [HK16] present parameterized families of exposed indecomposable maps, with explicit conditions for exposedness and the bi-spanning property. Collins, Hayden, and the second author [CHN16] employ random matrix and free probability techniques, constructing $k$-positive maps via free convolution and random matrix models. Marciniak and Rutkowski [MR17] introduce a geometric “merging” technique, systematically combining two positive maps with additional operators and functionals to produce new positive, often exposed, indecomposable maps. Müller-Hermes [MH18] uses symmetrization and tensor product techniques, particularly with Werner maps, to analyze decomposability under tensor powers and construct new non-decomposable maps. Siudzińska [Siu22] uses symmetric measurements (positive operator-valued measures) and operator bases to construct positive maps and indecomposable entanglement witnesses.
Other than or beyond the PPT criterion, the study of bipartite entanglement has also motivated the study of the notion of extendibility [DPS04, CKMR07, NOP09, Doh14, HNW17]. In particular, the notion of bosonic (or Bose-symmetric) extendibility plays a central role in the study of quantum entanglement, providing one of the most powerful and systematic separability criteria. Introduced in the seminal work of Doherty, Parrilo, and Spedalieri [DPS04], the hierarchy of $r$-extendible states forms an increasing sequence of convex sets that converges to the set of separable states. This completeness ensures that every entangled state eventually fails some level of the hierarchy. Moreover, each level can be efficiently tested using semidefinite programming (SDP), making the DPS hierarchy not only theoretically elegant but also computationally tractable.
A stronger version of this idea arises when one focuses on bipartite bosonic systems. In this setting, one may further require that the Bose-symmetric extensions simultaneously include both original subsystems, leading to what is called $r$-PPT bosonic exchangeability [CKMR07] or PPT extendible with respect to the complete graph $\mathcal{K}_{r}$ [ACG+23]. This refinement yields a sharper and still complete hierarchy of tests that interpolate between the PPT and separability conditions. Thanks to the monogamy of entanglement [Ter04], a state passing all these tests behaves as if it could be consistently extended to many identical copies while maintaining only a weak amount of entanglement across any bipartition. Despite the natural formulation of this criterion, explicitly constructing entangled states that satisfy such strong extendibility conditions remains a challenging problem, even for the simplest cases [DPS04].
@@
A powerful approach to understanding cones of maps and states is through duality. The cone of positive maps is dual to the cone of separable states, while the cone of decomposable maps is dual to the cone of PPT states [EK00, SSŻ09, Kye23]. Duality also connects map properties to cones of matrices. The cone of completely positive matrices $\mathsf{CP}_{n}$ [BSM03], consisting of matrices that can be factored as $XX^{\top}$ for some entrywise non-negative (rectangular) matrix $X$, has as its dual the cone of copositive matrices $\operatorname{\mathsf{COP}}_{n}$ [SMB21]. A real symmetric matrix $M$ is in $\operatorname{\mathsf{COP}}_{n}$ if $\langle x,Mx\rangle\geq 0$ for all non-negative vectors $x\in\mathbb{R}^{n}_{+}$. Just like the set of positive maps, the cone $\operatorname{\mathsf{COP}}_{n}$ is notoriously difficult to characterize, with membership being a co-NP-complete problem [MK87]. The subset of $\operatorname{\mathsf{COP}}_{n}$ comprising matrices that are a sum of a positive semidefinite and an entrywise non-negative matrix, denoted $\mathsf{SPN}_{n}$, is tractable and corresponds to the first level of a sum-of-squares (SOS) hierarchy that approximates $\operatorname{\mathsf{COP}}_{n}$. The strict containment $\operatorname{\mathsf{COP}}_{n}\supsetneq\mathsf{SPN}_{n}$ for $n\geq 5$, exemplified by the famous Horn matrix [SMB21], mirrors the distinction between positive and decomposable maps. Let us point out here that the use of copositive matrices for entanglement detection dates back to [MATS21], see also [Mar23, RPAR+25, MMRRP+25]. On the dual side, the connection between completely positive matrices and separability dates back to [Yu16, TAQ+18], see also [JM19, SN21, Sin21, Ber23, GNS25, Gul25]. Finally, the dual of (PPT-) bosonic extendibility hierarchy admits a natural interpretation in terms of sum-of-squares (SOS) representations of Hermitian forms [DPS04, FF21]. This duality connects the feasibility of an $r$-extendibility test to the existence of SOS certificates of block-positivity for corresponding entanglement witnesses. In this context, a fundamental open question was raised in [DPS04, Sec. VI], which we state informally,
Can the cone of entanglement witnesses (i.e. block-positive operators or elements of the dual of the separable set) be obtained as the union of the witnesses of these DPS levels?
We refer to Section 8.1 for further details. Any counterexample to this question could be used to detect entanglement of quantum states that are $r$-PPT bosonic extendible for arbitrarily high $r\geq 1$. While partial results are known (e.g., inclusion of its interior), a full resolution remains open, reflecting the deep interplay between convex geometry, operator algebras, and optimization in the characterization of quantum entanglement.
@@
On the other hand, a fruitful strategy for finding PPT entangled states and positive indecomposable maps is to impose symmetries on the states and maps under consideration [VW01, EW01, SN21, PJPY24]. A particularly rich and physically relevant class is that of diagonal unitary covariant (DUC) maps [Liu15, SN21], which commute with conjugations by diagonal unitary matrices. Such maps, and their close relatives the conjugate diagonal unitary covariant (CDUC) maps, have recently been the subject of intense study due to their structured algebraic properties [SN22, KMP18, SDN24]. Recent work established a fundamental isomorphism between the space of DUC/CDUC maps and the vector space of pairs of matrices $(A,B)$ sharing a common diagonal [JM19, SN21]. Within this framework, a cone of pairwise completely positive matrices $\operatorname{\mathsf{PCP}}_{n}$ was introduced, which was shown to correspond precisely to the class of entanglement breaking DUC/CDUC maps. This breakthrough provided a complete description of the strongest form of positivity for these maps. However, a corresponding characterization for general positivity was left as a significant open problem. This gap motivates the central questions of this paper:
@@
    Can these findings be further leveraged to construct families of entangled quantum states that nevertheless display separability-like properties, such as being PPT or possessing high-order symmetric extensions? More precisely, are there explicit examples of entangled bosonic states that are $\mathcal{K}_{r}$-PPT bosonic extendible? On the dual side, can we construct entanglement witnesses that admit no SOS certificate of any finite order?
@@
  • A bridge between the classical and pairwise cones: We establish a powerful constructive result, Theorem 4.1, which provides a canonical way to “lift” any matrix $M\in\operatorname{\mathsf{COP}}_{n}$ to a family of pairs in $\operatorname{\mathsf{PCOP}}_{n}$. This theorem enables the transfer results about copositive matrices into the domain of operator algebra and quantum information theory. These results, discussed in Section 4, and the results discussed in the next point (see Section 5) answer Q2.
  • Characterization of (in)decomposability: We extend our analysis to decomposability by introducing the cone of pairwise decomposable matrices, $\operatorname{\mathsf{PDEC}}_{n}$. We show this cone precisely characterizes decomposable DUC/CDUC maps and prove an analogous constructive theorem (Theorem 5.6) that links $\operatorname{\mathsf{PDEC}}_{n}$ to the tractable cone $\mathsf{SPN}_{n}$. In particular, for each matrix in $\operatorname{\mathsf{COP}}_{n}\setminus\mathsf{SPN}_{n}$, we construct an $(n^{2}-n)/2$-dimensional family of positive indecomposable linear maps (Corollary 5.9). Interestingly, we further show that when the underlying copositive matrix is extremal, the lifted family translates directly into a family of optimal entanglement witnesses, yielding an $(n^{2}-n)/2$-dimensional collection of “edge” objects (Corollary 5.10). The gap between $\operatorname{\mathsf{PCOP}}_{n}$ and $\operatorname{\mathsf{PDEC}}_{n}$ thus corresponds to the domain of indecomposable positive maps, and our results bridge this gap to the one between $\operatorname{\mathsf{COP}}_{n}$ and $\mathsf{SPN}_{n}$, which has been studied extensively in the literature. Markov-Choi maps (discussed in Section 6) correspond to LDUI maps having a special form. We characterize positivity, decomposability, and other properties of such maps in Proposition 6.1. In Propositions 6.3 and 6.6, we further provide new conditions for identifying positive indecomposable linear maps within this family. As a consequence, in Section 6.1 and Examples 6.4, 6.7, and 6.8, we present several examples of indecomposable Markov–Choi maps that significantly extend the known classes described in [Kye92, CKL92, Ha03].
@@
  • Connections to quantum state extendibility: In the final part of the paper, we connect our results to hierarchies of quantum state extendibility. We show in Proposition 8.4, Theorems 8.9 and 8.11 that the well-known sum-of-squares (SOS) hierarchies for copositive matrices ($\mathsf{K}_{n}^{(r)}$, see Section 7 and also [DKP02, Sch79]) correspond directly to witnesses for different levels of PPT-bosonic extendibility of quantum states, such as (mixture of) Dicke states. In particular, such connections allow us to construct a surprisingly wide family of entanglement witnesses which do not have SOS properties of any order (Theorem 8.5), thus providing counterexamples to the question dating back to [DPS04]. Furthermore, on the dual side, we completely characterize the PPT bosonic extendibility of Dicke states (with respect to the complete graph $\mathcal{K}_{r+2}$) in terms of the dual set $(\mathsf{K}_{n}^{(r)})^{\circ}$ (Theorem 8.16), which allows us to provide explicit examples of bosonic entangled states that are arbitrarily highly extendible whenever the local dimension $n$ is not less than $5$. This answers Q4. As a byproduct, we also prove in Corollary 8.18 the existence of $r$-partite entangled bosonic (Dicke) states of which are PPT with respect to all bipartitions.
The paper is organized as follows. Section 2 establishes our notation and reviews the necessary preliminaries on matrix cones and positive maps. Section 3 introduces the central object of study, the pairwise copositive cone $\operatorname{\mathsf{PCOP}}_{n}$, and proves its duality with $\operatorname{\mathsf{PCP}}_{n}$ as well as some other basic properties. Section 4 establishes the main constructive theorem linking the $\operatorname{\mathsf{PCOP}}_{n}$ cone to $\operatorname{\mathsf{COP}}_{n}$, hence providing a characterization of positivity for large classed of (C)DUC linear maps. Section 5 develops the analogous theory for the pairwise decomposable cone $\operatorname{\mathsf{PDEC}}_{n}$. In Section 6 we discuss Markov-Choi linear maps and their positivity properties. Section 7 applies this framework to define the graph-based maps and determines the exact positivity and decomposability thresholds; we also specialize this analysis to the case of strongly regular graphs. Finally, in Section 8 we put forward the remarkable connection between hierarchies of quantum state extendibility and hierarchies of sum-of-squares programs for membership in the $\operatorname{\mathsf{COP}}_{n}$ cone.